OpenStudy (anonymous):

how to write each expression with a single positive exponent? A) 4^4 x 4^6 x 4 B)8-^6 8-^5 C)m-^4 x m^4

OpenStudy (mertsj):

What is: \[x^4 \times x^6 \times x\]

OpenStudy (anonymous):

\[4^{4} \times 4^{6} \times 4\]

OpenStudy (mertsj):

If you will answer my question, I will help you.

OpenStudy (anonymous):

256 x 4096 x 4 = 4194304?

OpenStudy (mertsj):

\[x^4\times x^6\times x = x ^{11}\]

OpenStudy (mertsj):

The rule is that you add the exponents.

OpenStudy (mertsj):

So if you substitute 4 for x you get 4^11

OpenStudy (anonymous):

so would it be 4 x 4 x 4^11? or what

OpenStudy (mertsj):

\[4^4\times 4^6 \times 4^1 = 4^{4+6+1} = 4^{11}\]

OpenStudy (anonymous):

is that the final answer?